The local middle school had a concert to raise money. Tickets were $6.00 for adults and $4.00 for children. The committee figured there were 480 in attendance and that $2340.00 was raised. How many adults and how many children attended?
Answer by Ani_g(8)

Let the number of adults = x & the number of children = y
Then, x+y=480....(1)
& 6x+4y=2340....(2)
Now, by eq(1)x 4 => 4x+4y=1920....(3)
Next, by eq(2)-eq(3), we get, 2x=420 or x=210
Therefore, y=480-210=270.